Given the following term structure of risk-free interest rates today, what would you expect the interest rate to be on a one year bond four years in the future? Enter your answer as a percent without the “%.” Round your final answer to two decimals. Maturity in Years Interest Rate 1 4.00% 2 4.50% 3 4.75% 4 5.00% 5 6.00%
Answer #1

interest rates are as follow:

1 year spot rate 1r1 = 4%

2 year spot rate 1r2 = 4.5%

3 year spot rate 1r3 = 4.75%

4 year spot rate 1r4 = 5%

5 year spot rate 1r5 = 6%

to calculate, one year rate four years from now E(4r1)

So, E(4r1) = ((1+1r5)^5)/((1+1r4)^4) - 1 = (1.06^5)/(1.05^4) - 1 = 10.10%

So, one year rate four years from now is 10.10%